OpenStudy (anonymous):

2x^2+11x-21 what is the polynomial for this, AND HOW?? Plz & thx

OpenStudy (hba):

Use breaking the middle term.

OpenStudy (anonymous):

I still don't understand how to solve the problem

OpenStudy (hba):

Thinking of two digits which add up to 11 and when you multiply them they become -42

OpenStudy (hba):

@flyqueen I hope your question is correct and there is no sign problem ?

OpenStudy (anonymous):

I still don't understand can u solve it in steps for me plzzz and thank u for ur time

OpenStudy (hba):

I am off to sleep @sirm3d Please help.

OpenStudy (anonymous):

Good nite @hba

OpenStudy (sirm3d):

factoring problem?

OpenStudy (anonymous):

Yea

OpenStudy (sirm3d):

we'll try hba's method. can you think of two numbers (can be negative) that add up to 11 and whose product is -42? the product is negative so that means one of the numbers is positive and the other is negative.

OpenStudy (anonymous):

Umm I'm not sure what hba was talking about. So I'm kind of confused with thinking of a #

OpenStudy (sirm3d):

well, the 11 is the coefficient of the middle term, while -42 is the product of the coefficients of the first and last terms. suppose your number is 7 and (-6). 7(-6) = -42 but 7 + (-6) = 1 \(\neq\) 11. so 7 and (-6) are not the right pair.

OpenStudy (sirm3d):

can you think of a different number pair that add up to 11 and whose product is -42?

OpenStudy (sirm3d):

okay, i'll give the numbers, 14 and (-3). \[\large \begin{equation} 2x^2 +11x-21 \\ \quad = 2x^2+(14-3)x-21\\ \quad = 2x^2 + 14x - 3x - 21 \\ \quad = (2x^2+14x)+(-3x-21) \\\quad = 2x(x+7) -3(x+7) \\\quad = (2x-3)(x+7)\end{equation}\]

OpenStudy (anonymous):

@sirm3d Wouldn't slide and divide be easier? X^2+11x-42 (X+14)(x-3) (X+14/2)(x-3/2)(2) (X-7)(x-3)

OpenStudy (anonymous):

Sorry it's: (x+7)(2x-3)

OpenStudy (sirm3d):

it works for you, so that's all right. but others may not see the trick and get lost in the solution.

OpenStudy (anonymous):

Just a suggestion. Slide the coefficient of X and multiply it by the third term then divide the coefficient of x by that product.

OpenStudy (anonymous):

Then just do foil to check your answer